# Merging Rate of Opinions via Optimal Transport on Random Measures

### v2

@@ 2.2 Distances between random measures
The formula (4) makes sense for any ground metric space $(\mathbb{Y},d_{\mathbb{Y}})$, and we will use it with $(\mathbb{Y},d_{\mathbb{Y}})=(\mathcal{M}_{B}(\mathbb{X}),\mathrm{BL})$. Indeed, interpreting laws of random measures as probability distributions over $\mathcal{M}_{B}(\mathbb{X})$, it is possible to lift the distance $\mathrm{BL}$ into a distance over the space $\mathcal{P}(\mathcal{M}_{B}(\mathbb{X}))$. We call the resulting distance the “Wasserstein over Bounded Lipschitz” distance, denoted by $\mathcal{W}_{\mathrm{BL}}$. If we were to look at laws of random probability measures, that is, elements of $\mathcal{P}(\mathcal{P}(\mathbb{X}))$, then the analogue of the Wasserstein over Bounded Lipschitz distance would be the “Wasserstein over Wasserstein” distance, which has been proposed and studied in different contexts (Nguyen, 2016, Yurochkin et al., 2019, Dukler et al., 2019, Catalano and Lavenant, 2024).
@@
A random measure $\tilde{\mu}$ has finite mean if $\mathbb{E}(\tilde{\mu}(\mathbb{X}))<+\infty$. In this case its mean measure $\mathbb{E}(\tilde{\mu})$, defined by $\mathbb{E}(\tilde{\mu})(A)=\mathbb{E}(\tilde{\mu}(A))$, is an element of $\mathcal{M}_{B}(\mathbb{X})$.
@@
The quantity $\mathcal{W}_{\mathrm{BL}}$ defines a complete distance over the laws of random measures with finite mean. If $(\tilde{\mu}_{n})_{n}$, and $\tilde{\mu}$ are random measures with finite mean, then $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}_{n}),\mathcal{L}(\tilde{\mu}))$ converges to 00 as $n\to+\infty$ if and only if $(\tilde{\mu}_{n})_{n}$ converges in distribution to $\tilde{\mu}$ and $\mathbb{E}(\tilde{\mu}_{n}(\mathbb{X}))$ converges to $\mathbb{E}(\tilde{\mu}(\mathbb{X}))$.
@@ 3 Completely random measures and bounds via optimal transport on Lévy intensities
Recall that $\mathcal{M}(\mathbb{Y})$ is the set of locally finite non-negative measures, possibly with infinite mass. Mostly we consider $\mathbb{Y}=(0,+\infty)$ or $\mathbb{Y}=(0,+\infty)\times\mathbb{X}$. Here and after, =$\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}$ denotes equality in distribution.
@@
An important case is when $\rho_{x}$ does not depend on $x$. Typical densities used are $\rho(s)=\alpha s^{-1}\exp(-bs)$ or $\rho(s)=\alpha s^{-(1+\sigma)}\Gamma(1-\sigma)^{-1}\exp(-bs)$ corresponding to respectively a Gamma CRM and a generalized Gamma CRM, where $\alpha,b$ and $\sigma>0$ are parameters.
@@
$\iint_{(\varepsilon,+\infty)\times A}s\,\mathrm{d}\tilde{\mathcal{N}}(s,x)\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\sum_{i=0}^{N}s_{i}\,\delta_{x_{i}},$ (9)
@@
We move on to the definition of a distance at the level of Lévy intensities, which is actually related to the distance $\mathcal{W}_{\mathrm{BL}}$ defined in the previous section. The underlying idea is simple: if we can find a coupling between the Lévy intensities $\nu^{1},\nu^{2}$ and if $\tilde{\mu}^{i}\sim\textup{CRM}(\nu^{i})$ for $i=1,2$, then we have a coupling between $\tilde{\mu}^{1}$ and $\tilde{\mu}^{2}$, which means we should be able to find an upper bound of $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}),\mathcal{L}(\tilde{\mu}^{2}))$. The challenge is to understand how $\mathbb{E}(\mathrm{BL}(\tilde{\mu}^{1},\tilde{\mu}^{2}))$ relates to a quantity which depends only on the Lévy intensities.
This leads to the definition of a discrepancy between Lévy intensities $\mathrm{d}\nu(s,x)=\mathrm{d}\rho_{x}(s)\mathrm{d}P_{0}(x)$, which is ultimately justified by Theorem 4 below. We observe that the two components $\rho_{x}$ and $P_{0}$ play very different roles, as highlighted in Remark 1. Indeed, $P_{0}$ is a probability measure on $\mathbb{X}$ describing the law of the atoms, whereas $\rho_{x}$ is $P_{0}$-a.s. a measure with infinite mass on $(0,+\infty)$ that determines the conditional law of their corresponding jumps. We will thus preserve their different nature, in the spirit of the nested Wasserstein distance for probability measures, which was originally introduced by Pflug and Pichler (2012) and is often termed adapted Wasserstein distance in financial mathematics (Backhoff et al., 2022, Bartl et al., 2024). Specifically we impose that, if an atom $x$ is coupled to an atom $y$ on the space $\mathbb{X}$, then the whole measure $\rho_{x}$ is also coupled to $\rho_{y}$.
A key tool is a definition of an “extended” Wasserstein distance to compare measures of infinite mass $\rho_{x}$ and $\rho_{y}$ A (quadratic) Wasserstein distance between measures of different and potentially infinite mass has been proposed by Figalli and Gigli (2010) through the notion of extended coupling. We refer to Catalano et al. (2024) and Guillen et al. (2019) for a concise overview. In the present work, we will only consider measures on the half-real line $\mathbb{Y}=(0,+\infty)$, and look at Wasserstein distances of order $1$. We refer to Catalano et al. (2024, Proposition 4) for a link between the definition below and the previous works of Figalli and Gigli (2010) and Guillen et al. (2019).
@@
The extended Wasserstein distance (of order $1$) between two positive Borel measures $\rho^{1},\rho^{2}$ on $(0,+\infty)$ with finite first moments about 00 is defined as
@@
Following the aforementioned literature, and using the extended Wasserstein distance, we now present what we call the “adapated extended Wasserstein discrepancy”.
@@
The presence of the first moments $M_{1}(\nu^{1})$ and $M_{1}(\nu^{2})$ may seem counterintuitive, and prevents $\mathrm{AD}$ from being a distance. It is however necessary in order for the next theorem to hold. Note also that we use the truncated distance $d_{\mathbb{X},\text{t}}$ instead of the original distance $d_{\mathbb{X}}$. This is again because of the next theorem as it gives a sharper upper bound; see also (7) and the discussion related to it to understand why $d_{\mathbb{X},\text{t}}$ is a natural distance to consider. Thanks to the use of this truncated distance, note that there is no finiteness of any moment of $P_{0}^{i}$, $i=1,2$ in the definition of $\mathrm{AD}$. Actually, from Fubini $\int_{0}^{+\infty}U_{\rho}(s)\,\mathrm{d}s=\int_{0}^{+\infty}s\rho(s)\mathrm{d}s$ and thus $\mathcal{W}_{*}(\rho^{1}_{X},\rho^{2}_{X})\leq M_{1}(\nu^{1})+M_{2}(\nu^{2})$. It yields, together with the boundedness of $d_{\mathbb{X},\text{t}}$, that
@@
The adapted extended Wasserstein discrepancy $\mathrm{AD}$ is particularly convenient when at least one of the two Lévy intensities is homogeneous, i.e., $\nu^{1}=\rho^{1}\otimes P_{0}^{1}$. In such case, $\mathcal{W}_{*}(\rho_{X}^{1},\rho_{Y}^{2})^{p}=\mathcal{W}_{*}(\rho^{1},\rho_{Y}^{2})^{p}$ does not depend on the coupling and thus
@@
Let $\tilde{\mathcal{N}}^{1}$ and $\tilde{\mathcal{N}}^{2}$ be two Poisson random measures of intensities $\alpha_{i}P^{i}_{0}$, for $i=1,2$, where $\alpha_{i}>0$ and $P^{i}_{0}$ is a probability on $\mathbb{X}$. Then $\tilde{\mathcal{N}}^{i}$ is a CRM with Lévy intensity $\nu^{i}=\alpha_{i}\delta_{1}(\mathrm{d}s)\mathrm{d}P^{i}_{0}(x)$. In particular they are homogeneous CRMs with $\mathcal{W}_{*}(\rho^{1},\rho^{2})=|\alpha_{1}-\alpha_{2}|$. Thus putting together Theorem 4 and Example 1,
@@
A random measure $\tilde{\mu}$ on $\mathbb{X}$ is a Cox CRM or a doubly stochastic CRM if there exists a random element $\tilde{\nu}$ valued in $\mathcal{M}((0,+\infty)\times\mathbb{X})$ such that $\tilde{\mu}|\tilde{\nu}\sim\textup{CRM}(\tilde{\nu})$.
@@
In particular, if we restrict to $\tilde{\nu}^{1},\tilde{\nu}^{2}\in\mathcal{M}_{m}((0,+\infty)\times\mathbb{X})$ for some fixed m¿0, with $\mathcal{W}_{\mathrm{AD}}$ the Wasserstein distance over the $\mathrm{AD}$ metric in Definition 4, we have
@@ 4 Normalization, identifiability, and scaling
For the normalization $\tilde{\mu}/\tilde{\mu}(\mathbb{X})$ to be well-defined one needs $\tilde{\mu}(\mathbb{X})>0$ a.s., which at the level of the Lévy intensities corresponds to $\nu((0,+\infty)\times\mathbb{X})=+\infty$ (Regazzini et al., 2003). We will require a slightly stronger condition, namely that the CRM is infintely active. In terms of the canonical decomposition of Proposition 3, this assumption reads
for $\text{for }P_{0}\text{-a.e. }x,\qquad\rho_{x}(0,+\infty)=+\infty.$
This means that there will always be an infinite number of jumps in a set $A$ with $P_{0}(A)>0$, and thus the compound Poisson approximation (9) cannot be extended to $\epsilon=0$. In particular, it guarantees that for any set $A$ exactly one of the two alternatives holds: either $P_{0}(A)>0$, and in this case $\tilde{\mu}(A)>0$ a.s., or $P_{0}(A)=0$, which implies $\tilde{\mu}(A)=0$ a.s. This will be useful for our identifiability results.
@@ 4.1 Completely random measures
$\frac{\tilde{\mu}^{1}}{\tilde{\mu}^{1}(\mathbb{X})}\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\frac{\tilde{\mu}^{2}}{\tilde{\mu}^{2}(\mathbb{X})}\qquad\text{ if and only if }\qquad\tilde{\mu}^{1}\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\alpha\tilde{\mu}^{2}\text{ for some scalar }\alpha>0.$
@@
By Proposition 3 the mean measure $Q(\cdot)=\mathbb{E}(\tilde{\mu}(\cdot))$ of any scaled CRM with Lévy intensity $\mathrm{d}\nu(s,x)=\mathrm{d}\rho_{x}(s)\mathrm{d}P_{0}(x)$ is the probability measure $P_{0}$ and, for $P_{0}$-a.e. $x\in\mathbb{X}$,
@@
We highlight that, unlike normalized CRMs, scaled CRMs are CRMs. The following result provides a general way to find their Lévy intensities. We state it here for diffuse jumps components $\mathrm{d}\rho_{x}(s)=\rho_{x}(s)\,\mathrm{d}s$ for simplicity, though it can be extended to atomic jump components; see Lemma SM3.
@@
Thus we can regard $\tilde{\mu}_{\mathcal{S}}$ as a generalized gamma CRM that also accounts for a scale parameter.
When comparing normalized CRMs we should thus restrict their scaled version. Importantly, Proposition 5 and Theorem 4 guarantee that
@@
is a distance on on Lévy intensities of scaled CRMs that satisfies $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2}_{\mathcal{S}}))\leq\mathrm{AD}(\nu^{1}_{\mathcal{S}},\nu^{2}_{\mathcal{S}})$. Since Lévy intensities uniquely characterize the law of the CRM, the distance $\mathrm{AD}$ can be considered as a distance between CRMs: arguably a less canonical but more tractable one.

### v3

@@ 2.2 Distances between random measures
The definition in (4) makes sense for any ground metric space $(\mathbb{Y},d_{\mathbb{Y}})$, and we will use it with $(\mathbb{Y},d_{\mathbb{Y}})=(\mathcal{M}_{B}(\mathbb{X}),\mathrm{BL})$. Indeed, interpreting laws of random measures as probability distributions over $\mathcal{M}_{B}(\mathbb{X})$, it is possible to lift the distance $\mathrm{BL}$ into a distance over the space $\mathcal{P}(\mathcal{M}_{B}(\mathbb{X}))$. We call the resulting distance the “Wasserstein over Bounded Lipschitz” distance, denoted by $\mathcal{W}_{\mathrm{BL}}$. If we were to look at laws of random probability measures, that is, elements of $\mathcal{P}(\mathcal{P}(\mathbb{X}))$, then an analog of the Wasserstein over Bounded Lipschitz distance could be the “Wasserstein over Wasserstein” distance, which has been proposed and studied in different contexts (Nguyen, 2016, Yurochkin et al., 2019, Dukler et al., 2019, Catalano and Lavenant, 2024). We note that our Wasserstein over BL can be used on laws of random probabilities as well and requires weaker assumptions on first moments than the Wasserstein over Wasserstein distance.
@@
[…]
@@
The quantity $\mathcal{W}_{\mathrm{BL}}$ defines a complete distance over the laws of random measures with finite mean. If $(\tilde{\mu}_{n})_{n}$, and $\tilde{\mu}$ are random measures with finite mean, then $)\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}_{n}),\mathcal{L}(\tilde{\mu}$)) converges to 0 as $n\to+\infty$ if and only if $(\tilde{\mu}_{n})_{n}$ converges in distribution to $\tilde{\mu}$ and $\mathbb{E}(\tilde{\mu}_{n}(\mathbb{X}))$ converges to $\mathbb{E}(\tilde{\mu}(\mathbb{X}))$.
@@ 3 Completely random measures and bounds via optimal transport on Lévy intensities
Recall that $\mathcal{M}(\mathbb{Y})$ is the set of locally finite non-negative measures, possibly with infinite mass. Mostly we consider $\mathbb{Y}=(0,+\infty)$ or $\mathbb{Y}=(0,+\infty)\times\mathbb{X}$. Here and after, $\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}$ denotes equality in distribution.
@@
$\iint_{(\epsilon,+\infty)\times A}s\,\mathrm{d}\tilde{\mathcal{N}}(s,x)\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\sum_{i=0}^{N}S_{i}\,\delta_{X_{i}},$ (9)
@@
We move on to the definition of a distance at the level of Lévy intensities that is related to the distance $\mathcal{W}_{\mathrm{BL}}$ defined in the previous section. The underlying idea is simple: if we can find a coupling between the Lévy intensities $\nu^{1},\nu^{2}$ and if $\tilde{\mu}^{i}\sim\textup{CRM}(\nu^{i})$ for $i=1,2$, then we have a coupling between $\tilde{\mu}^{1}$ and $\tilde{\mu}^{2}$, which means we should be able to find an upper bound of $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}),\mathcal{L}(\tilde{\mu}^{2}))$. The challenge is to understand how $\mathbb{E}(\mathrm{BL}(\tilde{\mu}^{1},\tilde{\mu}^{2}))$ relates to a quantity that depends only on the Lévy intensities.
This leads to the definition of a discrepancy between Lévy intensities $\mathrm{d}\nu(s,x)=\mathrm{d}\rho_{x}(s)\mathrm{d}P_{0}(x)$, which is ultimately justified by Theorem 4 below. We observe that the two components $\rho_{x}$ and $P_{0}$ play very different roles, as highlighted in Remark 1. Indeed, $P_{0}$ is a probability measure on $\mathbb{X}$ describing the law of the atoms, whereas $\rho_{x}$ is $P_{0}$-a.s. a measure with infinite mass on $(0,+\infty)$ that determines the conditional law of their corresponding jumps. We will thus preserve their different nature, in the spirit of the nested Wasserstein distance for probability measures, which was originally introduced by Pflug and Pichler (2012) and is often termed adapted Wasserstein distance in financial mathematics (Backhoff et al., 2022, Bartl et al., 2024). Specifically, we impose that if an atom $x$ is coupled to an atom $y$ on the space $\mathbb{X}$, then the whole measure $\rho_{x}$ is also coupled to $\rho_{y}$.
A key tool is the definition of an “extended” Wasserstein distance to compare measures of infinite mass $\rho_{x}$ and $\rho_{y}$. A (quadratic) Wasserstein distance between measures of different and potentially infinite masses has been proposed by Figalli and Gigli (2010) through the notion of extended coupling. We refer to Catalano et al. (2024) and Guillen et al. (2019) for a concise overview. In the present work, we will only consider measures on the half-real line $\mathbb{Y}=(0,+\infty)$, and look at Wasserstein distances of order $1$. We refer to Catalano et al. (2024, Proposition 4) for a link between the definition below and the previous works of Figalli and Gigli (2010) and Guillen et al. (2019).
@@
The extended Wasserstein distance (of order $1$) between two positive Borel measures $\rho^{1},\rho^{2}$ on $(0,+\infty)$ with finite first moments about 0 is defined as
@@
Following the aforementioned literature, and using the extended Wasserstein distance, we now present what we call the “adapted extended Wasserstein discrepancy”.
@@
The presence of the first moments $M_{1}(\nu^{1})$ and $M_{1}(\nu^{2})$ may seem counterintuitive, and prevents $\mathrm{AD}$ from being a distance. It is however necessary in order for the next theorem to hold. Note also that we use the truncated distance $d_{\mathbb{X},\text{t}}$ instead of the original distance $d_{\mathbb{X}}$. This is again because of the next theorem as it gives a sharper upper bound; see also (7) and the discussion related to it to understand why $d_{\mathbb{X},\text{t}}$ is a natural distance to consider. Thanks to the use of this truncated distance, note that we do not need to require the finiteness of any moment of $P_{0}^{i}$ in the definition of $\mathrm{AD}$, for $i=1,2$. Moreover, $\int_{0}^{+\infty}U_{\rho}(s)\,\mathrm{d}s=\int_{0}^{+\infty}s\rho(s)\mathrm{d}s$ by Fubini’s theorem. Thus, $\mathcal{W}_{*}(\rho^{1}_{X},\rho^{2}_{X})\leq M_{1}(\nu^{1})+M_{2}(\nu^{2})$, which together with the boundedness of $d_{\mathbb{X},\text{t}}$ yields
@@
The adapted extended Wasserstein discrepancy $\mathrm{AD}$ is particularly convenient when at least one of the two Lévy intensities is homogeneous, i.e., $\nu^{1}=\rho^{1}\otimes P_{0}^{1}$ or, equivalently, $\rho_{X}^{1}$ does not depend on $X$ a.s., with $X\sim P_{0}$. Many of the most commonly used CRMs fall into this framework, such as Gamma CRMs, with $\rho(s)=\alpha s^{-1}\exp(-bs)$, or generalized gamma CRMs, with $\rho(s)=\alpha s^{-(1+\sigma)}\Gamma(1-\sigma)^{-1}\exp(-bs)$, where $\alpha,b>0$ and $0\leq\sigma<1$ are parameters. In such case, $\mathcal{W}_{*}(\rho_{X}^{1},\rho_{Y}^{2})^{p}=\mathcal{W}_{*}(\rho^{1},\rho_{Y}^{2})^{p}$ does not depend on the coupling and thus
@@
Let $\tilde{\mathcal{N}}^{1}$ and $\tilde{\mathcal{N}}^{2}$ be two Poisson random measures of intensities $\alpha_{i}P^{i}_{0}$, for $i=1,2$, where $\alpha_{i}>0$ and $P^{i}_{0}$ is a probability on $\mathbb{X}$. Then $\tilde{\mathcal{N}}^{i}$ is a CRM with Lévy intensity $\mathrm{d}\nu^{i}=\alpha_{i}\delta_{1}(\mathrm{d}s)\mathrm{d}P^{i}_{0}(x)$. In particular they are homogeneous CRMs with $\mathcal{W}_{*}(\rho^{1},\rho^{2})=|\alpha_{1}-\alpha_{2}|$. Thus, putting together Theorem 4 and Example 1,
@@
[…]
@@
In particular, if we restrict our attention to $\tilde{\nu}^{1},\tilde{\nu}^{2}\in\mathcal{M}_{m}((0,+\infty)\times\mathbb{X})$ for some fixed m¿0, with $\mathcal{W}_{\mathrm{AD}}$ the Wasserstein distance over the $\mathrm{AD}$ metric in Definition 4, we have
@@ 4 Normalization, identifiability, and scaling
For the normalization $\tilde{\mu}/\tilde{\mu}(\mathbb{X})$ to be well-defined one needs $\tilde{\mu}(\mathbb{X})>0$ a.s., which at the level of the Lévy intensities corresponds to $\nu((0,+\infty)\times\mathbb{X})=+\infty$ (Regazzini et al., 2003). We will require a slightly stronger condition, namely that the CRM is infinitely active. In terms of the canonical decomposition of Proposition 3, this assumption reads
$\rho_{X}(0,+\infty)=+\infty.$
almost surely, where $X\sim P_{0}$. This means that there will always be an infinite number of jumps in a set $A$ with $P_{0}(A)>0$, and thus the compound Poisson approximation (9) cannot be extended to $\epsilon=0$. In particular, it guarantees that for any set $A$ exactly one of the two alternatives holds: either $P_{0}(A)>0$, and in this case $\tilde{\mu}(A)>0$ a.s., or $P_{0}(A)=0$, which implies $\tilde{\mu}(A)=0$ a.s. This will be useful for our identifiability results.
@@ 4.1 Completely random measures
[…]
@@
[…]
@@
almost surely, where $X\sim P_{0}$. We highlight that, unlike normalized CRMs, scaled CRMs are CRMs. The following result provides a general way to find their Lévy intensities. We state it here for diffuse jumps components $\mathrm{d}\rho_{x}(s)=\rho_{x}(s)\,\mathrm{d}s$ for simplicity, though it can be extended to atomic jump components; see Lemma SM3 in the Supplementary Material (Catalano and Lavenant, 2025).
@@
Thus, we can regard $\tilde{\mu}_{\mathcal{S}}$ as a generalized gamma CRM that also accounts for a scale parameter.
When comparing normalized CRMs we should thus restrict our attention to their scaled version. Importantly, Proposition 5 and Theorem 4 guarantee that
@@
is a distance on Lévy intensities of scaled CRMs that satisfies $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2}_{\mathcal{S}}))\leq\mathrm{AD}(\nu^{1}_{\mathcal{S}},\nu^{2}_{\mathcal{S}})$. Since Lévy intensities uniquely characterize the law of CRMs, the adapted extended Wasserstein distance $\mathrm{AD}$ can be considered as a distance on CRMs: arguably a less canonical but more tractable one.